OpenStudy (anonymous):

Find the difference quotient f(x+h)-f(x)/h, where h doesn't equal 0 , for the function below: f(x)=3x^2-9 Simplify your answer as much as possible. I came up with 6x+3h, am I doing this right?

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h }, where (h) \neq 0\]

OpenStudy (tkhunny):

Okay, but you have not completed the exercise. Now, find the limit as h begins to vanish.

OpenStudy (anonymous):

how do I do that?

OpenStudy (tkhunny):

You ponder it. h = 10 ==> 6x + 30 h = 1 ==> 6x + 3 h = 0.1 ==> 6x + .3 h = 0.01 ==> 6x + .03 h = 0.001 ==> 6x + .003 h = 0.0001 ==> 6x + .0003 h = 0.00001 ==> 6x + .00003 h = 0.000001 ==> 6x + .000003 What's going on?

OpenStudy (anonymous):

the answer is 3, right? thank you for helping me

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@tkhunny, it doesn't look like finding the limit is required; just the difference quotient. @AwesomeAries, you're given \(f(x)=3x^2-9\), so \[\begin{align*}\frac{f(x+h)-f(x)}{h}&=\frac{3(x+h)^2-9-\left(3x^2-9\right)}{h}\\ &=\frac{3\left(x^2+2xh+h^2\right)-9-3x^2+9}{h}\\ &=\frac{3x^2+6xh+3h^2-9-3x^2+9}{h}\\ &=\frac{6xh+3h^2}{h}\\ &=6x+3h \end{align*}\] So yes, your answer is correct.

OpenStudy (tkhunny):

Fair enough.

OpenStudy (anonymous):

thank you both